Oxford Materials IT Committee Meeting

MICHAELMAS TERM 2009
A meeting of the Department of Materials IT Committee was held at 11am on Tuesday 27th October 2009 in the HR Meeting Room.

AGENDA

  • 1 Apologies for absence and Membership of the Committee
  • 2 Matters arising from the Minutes of the Meeting of the last meeting
  • 3 IT/AV – Matters for Discussion (attached)
  • 4 IT/AV – Head of IT Report (attached)
  • 5 IT/AV expenditure - matters for report (PJW)
  • 6 AOB
  • 7 Date and time of next meeting: Tuesday 2nd February 2010 at 11am in HR Meeting Room

ITEM 2. MATTERS ARISING FROM THE MEETING

See minutes below.

ITEM 3: IT/AV MATTERS FOR DISCUSSION

  • Windows 7
  • Labview

ITEM 4: IT/AV HEAD OF IT REPORT

The report below summarises activities since the last meeting of ITC, as categorised by topic. The report is intended for record rather than for discussion at ITC meeting although queries may arise.

Admin Network

Software upgrades associated with Financials upgrade.

Teaching Laboratory

No significant changes. No annual roll (again) due to budget constraints. Microscope desktops are struggling and the laptops are old, but there is not budget allocated for replacement.

Lecture theatres / meeting rooms

All rooms checked before term. Portable audio-visual equipment is in Rex-Richards meeting room and a quote is in progress for fixed equipment.

Digital Print-room

Print usage levels and toner costs are high and increasing. Note that the cost of colour toners increased by 30% in October (4p->6p/page). The PC’s are soon to be replaced.

Library

The self-issue computer monitor was changed to flat screen due to room refurbishments and one of the desktop monitors was replaced due to hardware failure. Both PC’s in the library are slow and need rolling replacements.

MML

A 192 core cluster was specified and installed in Rex Richards. The cluster Slater was reinstalled. All the MML desktops are being reinstalled. The main file-storage array failed twice and was decommissioned.

EM

The EM firewall is being decommissioned and all machines moved back onto the main Materials network. The Bewapass door-entry system is scheduled for replacement (no timescale). Tracey Finn replaced Clive Downing but has had little interaction with IT staff.

QIPIRC

The qipirc website was migrated to a different server and Vasti Gregory received some training. Several laptops were disposed to departing researchers. Several other older laptops were reinstalled for reuse.

Website

The department website now has database driven homepages and project listings. Further developments are in progress (slowly due to time constraints).

The new weblearn site went live in June 2009. The old weblearn site has been decommissioned.

https://weblearn.ox.ac.uk/portal/hierarchy/mpls/materials

All CMS websites were upgraded to fix security bugs.

Publications

Researches in Progress 2009 document has been started.

The printed mugshots have been updated for the new academic year.

The handbooks were produced for the new academic year. 585 handbooks with 35,500 printed sides costing £1500 in toner plus two weeks of fulltime.

Network

The Materials subnet went offline for 1hr on 20th October, and then for 51hrs on 24-26th due to a hardware failure of the FroDo UPS. Problems with network instability in Parks Road were resolved by moving the feed back onto internal fibres.

Security

Several users received help removing viruses.

Disposal

Disposal of old equipment is still waiting to be done.

Software

Security issues with Java, Adobe, TSM software are being addressed department-wide.

Disaster Recovery and Backup

The MML RAID failure was recovered from TSM rather than Amanda.

New device registrations

The new device management reporting shows 35 devices in October, 31 devices in September.


MINUTES

A meeting of the Department of Materials IT Committee was held at 11am on Tuesday 27th October 2009 in the HR Meeting Room.
The following people attended: APD, PJW, CA, JDM, RT, SCB.

1. Apologies for Absence and Membership of the Committee

a) Apologies received from AMC who was troubleshooting Materials0 and KS who was off.

2. Matters arising from the minutes

a) IT disaster recovery plans (Minute 2a)
There is still a need for disaster recovery documentation to ensure that there is adequate provision and there is a clear structure for achieving efficient disaster recovery and ensuring business continuity.

Action: PJW

b) Firewall exceptions (Minute 2d)
It was agreed that the department firewall policy should not allow blanket exceptions. It was agreed that PJW and APD should discuss this with Nicole Grobert. It was agreed that a webpage should be created in order to advertise the department firewall policy and rule sets. APD has made initial contact with NG but discussion was postponed due to NG skiing injury.

Action: APD/PJW

c) Update on Groupware project (Minute 2e)
IT staff migrated to Nexus in August and all users (except 15 users scheduled for November) were migrated in September with only three failures. GDWS had the most problem with email migration. In order for users to reap the benefits of using an exchange server IT staff will need to promote the use of Outlook2007 which is part of Office2007.

The Core User Directory project at http://www.ict.ox.ac.uk/odit/projects has completed a pilot project and produced a report. The CUD project is needing high level support to define policy upon authority and move to becoming a central service. Alex attended a presentation at the ITSS conference on 15th July.

d) Podcasting (Minute 2f)
PJW and SCB have met and discussed creating a Materials area on the Oxford iTunes site. The plan is to record 2-4 podcasts of recent prize-winning student presentations. Project was not taken forward as a summer project due to time constraints.

Action: PJW/SCB

PJW has done the initial conversion from VHS to AVI of all the video tapes located in the Teaching Laboratory. SGR has edited clips necessary for his lectures. PJW is to review and edit the recordings and link files to the Video Catalogue.

Action: PJW

e) WiFi at Begbroke (Minute 6d)
Wireless site-surveys have been undertaken across the department and so far three new OWL/EDUROAM access points have been installed (BR committee room, BR lecture theatre, HR library). The schedule for completing the migration from OWL-VPN/OWL-VISITOR to OWL/EDUROAM is dependant upon getting the remaining access points from OUCS.

Action: PJW/KS

f) Central File Storage (Minute 6e)
IT staff have been promoting usage of Materials0 but the summer audit of group IT activities did not occur so there has not been a systematic approach to all research groups.

AMC and PJW have modified the IT support web-pages to give clearer information about how to achieve common tasks. e.g. fileserver usage, printer installation. Further suggestions welcome.

g) Slow speed of computers in the Digital Printroom (Minute 2o)
Both computers are still waiting to be upgraded to provide a better facility for users. Budget has been allocated.

Action: AMC/CA

h) Weblearn (Minute 3a)
The new weblearn site went live in June 2009. The old weblearn site has been decommissioned. https://weblearn.ox.ac.uk/portal/hierarchy/mpls/materials

i) Finance update (Minute 5)
PJW circulated the end-of-year summary of expenditure.

3 Matters for Discussion

a) Windows 7 Support
Windows7 was officially released in October 2009. The existing policy on Windows from MT2008 is:

  • department machines should remain using Windows XP unless a clear justification for using Vista (only Business/Enterprise/Ultimate) is put forward by the user and accepted by IT staff
  • personally owned machines can use Vista but users should be notified that support from IT staff for problems with Vista will be low priority.

CA has been running Windows7 Beta for 6 months and confirmed that it seems a significant improvement over Windows Vista and is stable. Some core software is not yet available for Windows 7 (TSM,VPN) but overall IT staff are keen to move forward to being able to support Windows7 and will carry out some necessary systems developments.

Action: AMC/CA

b) Labview site-license
JDM presented a summary of discussions about Materials usage of Labview. (see below)

It was decided that current usage does not justify the £2875/yr contribution for site license however JDM and PJW should approach John MacAllister in Physics to see whether it is possible to negotiate a lower contribution in order to increase usage around the department.

Action: JDM/PJW

c) Mathematica site-license
Due to the Labview discussion, there was also mention of the Physics trying to raise interest in a site-license for undergraduate use of Mathematica. PJW noted that there was currently no undergraduate usage of Mathematica and therefore no motivation to pay ~£400/yr in addition to the existing £550/yr for research provision.

However Simon Benjamin (lecturer 2nd year Maths) indicated that he would be keen to incorporate some Mathematica into his lecture course if we did make Mathematica available to undergraduates. PJW and SCB to contact Ian MacAllister again to discuss.

Action: SCB/PJW

d) Review of Summer Tasks
Before the summer the IT support team have identified a list of projects and prioritized them in terms of whether they must be completed before the new academic year, whether they should be completed or whether they could be done if time permits. Below summarises progress towards that list.

Must do:

  • Nexus/Groupware (migration, user training, roll-out Office2007):
    Migration has occurred for almost all users. Users were provided support on request but not coordinated training has occurred. Office2007 is promoted but has not been rolled-out. (95% done)
  • Vista printing support (iPrint drivers):
    iPrint now supports Vista and also Mac and Linux clients. (100% done)
  • Servers (software upgrades to Mateirals0 and others):
    Numerous software upgrades were done (Netware, eDir, FilesystemFactory, Zenworks, Patches) but Materials0 remains in need of upgrade/replacement. (80% done)
  • Virtualization (planning, costing, implementation):
    Virtualization remains in use only for management systems rather than user systems. (10% done)
  • Device Audit / Security Audit (start rolling program across whole department):
    A new system for monitoring the status of devices on our network was implemented by Alex but a full device and security audit was not started. (10% done)
  • Website Intranet (personal pages, projects, publications):
    Personal pages and all research project and projects available are now database driven, and a web-interface for updating the data has been written but it not yet sufficiently robust for users to been given full access. (80% done)
  • Department Publications (printing/binding handbooks):
    The handbooks were produced for the new academic year. 585 handbooks with 35,500 printed sides costing £1500 in toner plus two weeks of fulltime. (100% done)
  • MML tasks (new cluster, reinstall old cluster, upgrade desktop OS):
    New cluster Kittel was specified and installed in Rex Richards. The cluster Slater was reinstalled. All the MML desktops are being reinstalled. (90% done)
  • IT Task Management (Request Tracker, Budgeting):
    Request Tracker was installed but has not been used in production. RT needs to be revisited when time permits. IT finances were reviewed for the end of the financial year. (50% done)

Should do:

  • Vista/Windows7 (increased support capabilities):
    We have rolled-out several computers with Vista 64bit but the installation and configuration requires significantly more intervention that rolling-out XP.
  • Weblearn (promote usage with AOT):
    The new weblearn site went live in June 2009. The old weblearn site has been decommissioned. https://weblearn.ox.ac.uk/portal/hierarchy/mpls/materials
  • VOIP@Begbroke (continue migrating Siemens to Cisco VOIP):
    No further progress on VOIP over the summer due to time constraints.
  • OWL Phase 2 (schedule for implementation, also Begbroke) (Minute 6d):
    Wireless site-surveys have been undertaken across the department and so far three new OWL/EDUROAM access points have been installed (BR committee room, BR lecture theatre, HR library). The schedule for completing the migration from OWL-VPN/OWL-VISITOR to OWL/EDUROAM is dependant upon getting the remaining access points from OUCS.
  • Identity Manager (develop framework, link to CUD project):
    No progress over the summer.
  • Server hardware upgrades (WSUS, Altiris):
    No progress over the summer.
  • Door entry systems (merger with EngSci?, plans for EM doors):
    No progress over the summer.
  • Subversion (configuration management for IT, software for EM and MML):
    A subversion server has bee implemented and some users added (Giustino, Lin). This server needs to be documented and advertised as a departmental service.
  • Electronic Archive (review questionnaire, develop policy):
    No progress over the summer.

Could do:

  • LowC ICT (roll-out WOL, user training, device monitoring):
    No progress over the summer apart from attending May workshop and July presentation.
  • Podcasting/VideoLibrary:
    No progress over the summer.

4. Head of IT Report

A written report was circulated with the agenda. (see above)

5. IT/AV Expenditure

PJW presented a summary of expenditure before year end on 31st July 2009. Total consumables expenditure was around £36.5k (within the £37k budget).

PJW reminded the committee that this year’s budget has been cut by 20% to £30k. Summary expenditure to end October is already £15k but most of that will be charged to research grants.

PJW noted that HP toner prices increased by 50% at the end of September. Given that last year more than £5k was budgeted for toners for the HPCLJ5550dn printers, this is potentially going to have a big impact upon this year’s already tight IT budget. APD requested PJW check with central purchasing, but PJW confirmed that changing suppliers still meant that the prices had increased by >20%. XMA seems no-longer to be the most cost-effective supplier of print consumables.

6. AOB

a) Committee membership
APD noted that Richard Turner’s term of office had expired. PJW recommended re-co-opting RT onto the committee as Begbroke representative.

Action: PJW

7. Date of Next Meeting

The next meeting will take place on Tuesday 2nd February 2010 at 11am in the HR Meeting Room (ETB and HR both unavailable at 2pm)

The meeting closed at 12.30pm